# Geometry of exactness of moment-SOS relaxations for polynomial optimization

### v3

@@ Abstract
The moment-SOS (sum of squares) hierarchy is a powerful approach for solving globally non-convex polynomial optimization problems (POPs) at the price of solving a family of convex semidefinite optimization problems (called moment-SOS relaxations) of increasing size, controlled by an integer, the relaxation order. We say that a relaxation of a given order is exact if solving the relaxation actually solves the POP globally. In this note, we study the geometry of the exactness cone, defined as the set of polynomial objective functions for which the relaxation is exact. Generalizing previous foundational work on quadratic optimization on real varieties, we prove by elementary arguments that the exactness cones are unions of semidefinite representable cones monotonically embedded for increasing relaxation order.
@@ 1 Solving POPs with the moment-SOS hierarchy
Given a compact semialgebraic set ${\mathscr{X}}\in\mathbb{R}^{n}$ and a polynomial $f$ in the vector space $\mathbb{R}[{\bf x}]_{d}$ of polynomials of ${\bf x}\in\mathbb{R}^{n}$ of degree up to $d$, consider the polynomial optimization problem (POP)
@@
If ${\mathscr{X}}:=\{{\bf x}\in\mathbb{R}^{n}:g_{i}({\bf x})\geq 0,i=1,\ldots,m\}$ is basic semialgebraic, defined by a given polynomial vector ${\bf g}=(g_{i})_{i=1,\ldots,m}$, then ${\mathscr{P}}({\mathscr{X}})_{d}$ can be approximated with other convex cones, the truncated quadratic modules
@@
Note that if a polynomial equation enters the definition of ${\mathscr{X}}$, i.e. $g_{i}({\bf x})=0$ instead of $g_{i}({\bf x})\geq 0$ for some $i=1,\ldots,m$, then the corresponding weight $s_{i}\in\mathbb{R}[{\bf x}]$ in the quadratic module ${\mathscr{Q}}({\bf g})^{r}_{d}$ is not constrained in sign, while satisfying $s_{i}g_{i}\in\mathbb{R}[{\bf x}]_{2r}$. This is consistent with the fact that two inequalities of opposite signs are equivalent to an equation. Without loss of generality, and for notational conciseness, in this note we use only inequalities.
@@
Assumption 1.
For large enough $R\in\mathbb{R}$ and $r\in\mathbb{N}$ it holds $R^{2}-\sum_{k=1}^{n}x^{2}_{i}\in{\mathscr{Q}}({\bf g})^{r}_{2}$.
Since ${\mathscr{X}}$ is bounded, it is always possible to add a redundant quadratic constraint $R^{2}-\sum_{k=1}^{n}x^{2}_{i}\geq 0$ to the description of ${\mathscr{X}}$. So Assumption 1 is without loss of generality. Then it follows from [6] that in (3) the primal is attained (i.e. the infimum is a minimum), there is no duality gap (i.e. the infimum equals the supremum) and the relaxed values are monotonically converging lower bounds on the value:
$\text{sos}(f)^{r}=\text{mom}(f)^{r}\leq\text{sos}(f)^{r+1}=\text{mom}(f)^{r+1}\leq\text{sos}(f)^{\infty}=\text{mom}(f)^{\infty}=v(f).$
Moreover if ${\mathscr{X}}$ has an interior point, then the dual is attained (i.e. the supremum is a maximum). If ${\mathscr{X}}$ does not have an interior point, e.g. if it is a low-dimensional algebraic variety, then additional algebraic or geometric conditions are required for the dual to be attained, cf. [2, 1, 8].
@@ 2 Exactness cone
${\mathscr{F}}({\bf g})^{r}_{d}:=\{f\in\mathbb{R}[{\bf x}]_{d}:\ell_{\hat{{\bf y}}}(f)=\hat{v}=v(f),\>\hat{{\bf y}}\in{\mathscr{R}}({\bf g})^{r}_{d},\>f-\hat{v}\in{\mathscr{Q}}({\bf g})^{r}_{d}\}.$
@@
Our analysis is elementary. It is inspired by the foundational work [3] which focused on the particular case of the Shor relaxation ($r=1$) with $f$ linear ($d=1$) or quadratic ($d=2$), and ${\mathscr{X}}$ a real algebraic variety defined by quadratic equations. We believe that our contribution consists of considerably simplifying and extending this analysis to higher order relaxations of general semialgebraic sets.
@@ 3 Main result
The exactness cone at relaxation order $r$ is given by
@@
Cone ${\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}$ is semidefinite representable since it is the projection of linear sections of the SOS cone $\Sigma[{\bf x}]$, which is itself semidefinite representable. $\Box$
@@
The exactness cone is semialgebraic since $f$ belongs to ${\mathscr{F}}({\bf g})^{r}_{d}$ whenever 1) $f$ belongs to ${\mathscr{Q}}({\bf g})^{r}_{d}$, a semidefinite representable hence semialgebraic cone, and 2) there exists $\hat{{\bf x}}$ in ${\mathscr{X}}$, a semialgebraic set, such that $f$ vanishes at $\hat{{\bf x}}$.
@@
It follows from the proof of Theorem 1 that it is enough to restrict the union (5) to points $\hat{{\bf x}}\in{\mathscr{X}}$ which are optimal for some objective function $f$. For example if $d=1$, the union can be restricted to the set of extreme points of the convex hull of ${\mathscr{X}}$. In general however it is not easy to describe explicitly these subsets of ${\mathscr{X}}$.
@@
[…]
@@ 4 Geometry of exactness cones
The exactness cones are monotonically embedded for increasing relaxation order:
@@
Proof: The inclusion $R[{\bf x}]_{0}\subset{\mathscr{F}}({\bf g})^{r}_{d}$ follows from the translation invariance of the value: $v(f+a)=v(f)+a$ for all $a\in\mathbb{R}$. The inclusion ${\mathscr{F}}({\bf g})^{r}_{d}\subset{\mathscr{F}}({\bf g})^{r+1}_{d}$ follows from the inclusion relations (2). Finally, the identity $\overline{{\mathscr{F}}({\bf g})^{\infty}_{d}}=\mathbb{R}[{\bf x}]_{d}$ follows from Putinar’s Positivstellensatz – see e.g. [7, Thm. 2.15] – which states that under Assumption 1 every strictly positive polynomial of degree $d$ on ${\mathscr{X}}$ belongs to ${\mathscr{Q}}({\bf g})^{r}_{d}$ for sufficiently large $r\in\mathbb{N}$, i.e. $\overline{{\mathscr{Q}}({\bf g})^{\infty}_{d}}={\mathscr{P}}({\mathscr{X}})_{d}$. $\Box$
@@
Note that in Lemma 1 the closure is required for asymptotic exactness of all degree $d$ polynomial objective functions, i.e. $\overline{{\mathscr{F}}({\bf g})^{\infty}_{d}}=\mathbb{R}[{\bf x}]_{d}$. A classical example is $n=m=1$, $g_{1}(x)=-x^{2}$ for which ${\mathscr{X}}=\{0\}$ and Assumption 1 is readily satisfied. Whereras $f(x)=\pm x\notin{\mathscr{Q}}({\bf g})^{r}_{1}$ for finite $r\in\mathbb{N}$, it holds that $f(x)+\varepsilon=\frac{\varepsilon}{2}+\frac{\varepsilon}{2}(1\pm\frac{x}{\varepsilon})^{2}-\frac{x^{2}}{2\varepsilon}+\in{\mathscr{Q}}({\bf g})^{1}_{1}$ for every $\varepsilon>0$, see [1, Ex. 1.3.4] or [8, §2.5.2]. For this example ${\mathscr{F}}({\bf g})^{r}_{1}=\mathbb{R}[{\bf x}]_{0}$ for all finite $r$, and $\overline{{\mathscr{F}}({\bf g})^{\infty}_{1}}=\mathbb{R}[{\bf x}]_{1}$.
Given $\hat{{\bf x}}\in{\mathscr{X}}$, let ${\bf y}_{\hat{{\bf x}}}$ be the Dirac vector at $\hat{{\bf x}}$, i.e. such that $\ell_{{\bf y}_{\hat{{\bf x}}}}(f)=f(\hat{{\bf x}})$ for every $f\in\mathbb{R}[{\bf x}]_{d}$. Given a convex set ${\mathscr{Y}}\in\mathbb{R}^{N}$, let ${\mathscr{N}}_{{\mathscr{Y}}}({{\bf y}})$ denote the normal cone to ${\mathscr{Y}}$ at a point ${{\bf y}}\in{\mathscr{Y}}$. Recall that $f\in{\mathscr{N}}_{{\mathscr{Y}}}({{\bf y}})$ if and only if ${{\bf y}}\in\mathrm{argmax}_{{\bf y}\in{\mathscr{Y}}}\ell_{{\bf y}}(f)$, where $\ell_{{\bf y}}(f)$ is a linear functional on $\mathbb{R}^{N}$. With these notations, we have the following geometric counterpart to Theorem 1.
Lemma 2.
Up to the sign, the spectrahedral cones of Theorem 1 are normal cones to the moment relaxation:
${\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}=-{\mathscr{N}}_{{\mathscr{R}}({\bf g})^{r}_{d}}({\bf y}_{\hat{{\bf x}}})$
@@
Proof: First observe that extreme points of the moment cone ${\mathscr{M}}({\mathscr{X}})_{d}$ are Dirac vectors of ${\mathscr{X}}$, cf. e.g. [8, Thm. 8.1.1]. Dirac vectors which are also extreme points of the moment relaxation ${{\mathscr{R}}({\bf g})^{r}_{d}}$ are of the form ${{\bf y}}_{\hat{{\bf x}}}$ for some optimal $\hat{{\bf x}}\in{\mathscr{X}}$ for POP (1).
Therefore, if $\hat{{\bf x}}$ is optimal then ${\bf y}_{\hat{{\bf x}}}\in\mathrm{argmin}_{{\bf y}\in{\mathscr{R}}({\bf g})^{r}_{d}}\ell_{{\bf y}_{\hat{{\bf x}}}}(f^{0})$ with $f^{0}:=f-v(f)\in-{\mathscr{N}}_{{\mathscr{R}}({\bf g})^{r}_{d}}({\bf y}_{\hat{{\bf x}}})$, $\ell_{{\bf y}_{\hat{{\bf x}}}}(f^{0})=f^{0}(\hat{{\bf x}})=0$. Since ${\mathscr{Q}}({\bf g})^{r}_{d}$ is dual to ${\mathscr{R}}({\bf g})^{r}_{d}$, it follows that $f^{0}\in{\mathscr{Q}}({\bf g})^{r}_{d}$ and hence $f^{0}\in{\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}$.
If $\hat{{\bf x}}\in{\mathscr{X}}$ is not optimal, then the Dirac vector ${\bf y}_{\hat{{\bf x}}}$ cannot be on the boundary of the moment relaxation ${{\mathscr{R}}({\bf g})^{r}_{d}}$, so the normal cone ${\mathscr{N}}_{{\mathscr{R}}({\bf g})^{r}_{d}}({\bf y}_{\hat{{\bf x}}})$ at this point is zero, which trivially belongs to ${\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}$. $\Box$
Lemma 2 is a generalization of [3, Prop. 4.7] which considers only the case $d\leq 2$ and the Shor relaxation (i.e. $r=1$) for real varieties ${\mathscr{X}}$ generated by quadratic polynomials.
It is of interest to know whether we have exactness at relaxation order $r$ for all objective functions. Obviously, a necessary and sufficient condition is that ${\mathscr{R}}({\bf g})^{r}_{d}={\mathscr{M}}({\mathscr{X}})_{d}$, i.e. all vectors of the moment relaxation are convex combinations of Dirac vectors of ${\mathscr{X}}$. If $d=1$ the condition becomes ${\mathscr{R}}({\bf g})^{r}_{1}=\mathbb{R}_{+}\times\mathrm{conv}\>{\mathscr{X}}$, the Cartesian product of the positive real line with the convex hull of ${\mathscr{X}}$.

### v4

@@ Abstract
The moment-SOS (sum of squares) hierarchy is a powerful approach for solving approximately and globally non-convex polynomial optimization problems (POPs) at the price of solving a family of convex semidefinite optimization problems (called moment-SOS relaxations) of increasing size, controlled by an integer, the relaxation order. We say that a relaxation of a given order is exact if solving the relaxation actually solves the POP globally. In this note, we study the geometry of the exactness cone, defined as the set of polynomial objective functions for which the relaxation is exact. Generalizing previous foundational work on quadratic optimization on real varieties, we prove by elementary arguments that the exactness cones are unions of semidefinite representable cones monotonically embedded for increasing relaxation order.
@@ 1 Solving POPs with the moment-SOS hierarchy
Given a compact semialgebraic set ${\mathscr{X}}{\>\subset\>}\mathbb{R}^{n}$ and a polynomial $f$ in the vector space $\mathbb{R}[{\bf x}]_{d}$ of polynomials in ${\bf x}\in\mathbb{R}^{n}$ of degree up to $d$, consider the polynomial optimization problem (POP)
@@
If ${\mathscr{X}}:=\{{\bf x}\in\mathbb{R}^{n}:g_{i}({\bf x})\geq 0,i=1,\ldots,m\}$ is basic semialgebraic, defined by a given polynomial vector ${\bf g}=(g_{i})_{i=1,\ldots,m}$, then ${\mathscr{P}}({\mathscr{X}})_{d}$ can be approximated from inside with other convex cones, the truncated quadratic modules
@@
Since set ${\mathscr{X}}$ is bounded, it is always possible to add a redundant ball constraint to its description. So the following assumption is without loss of generality.
Assumption 1.
Let $R>0$ be the radius of an Euclidean ball including ${\mathscr{X}}$. It holds $R^{2}-\sum_{k=1}^{n}x^{2}_{i}\in{\mathscr{Q}}({\bf g})^{r}_{2}$ for all $r\in\mathbb{N}$.
@@
Note that if a polynomial equation enters the definition of ${\mathscr{X}}$, i.e. $g_{i}({\bf x})=0$ instead of $g_{i}({\bf x})\geq 0$ for some $i=1,\ldots,m$, then the corresponding weight $s_{i}\in\mathbb{R}[{\bf x}]$ in the quadratic module ${\mathscr{Q}}({\bf g})^{r}_{d}$ is not constrained in sign, while satisfying $s_{i}g_{i}\in\mathbb{R}[{\bf x}]_{2r}$. Recalling the identity $s_{i}=((s_{i}+1)/2)^{2}-((s_{i}-1)/2)^{2}$, every polynomial of $\mathbb{R}[{\bf x}]$ can be written as the difference of two polynomials of $\Sigma[{\bf x}]$. This is consistent with the fact that two inequalities of opposite signs are equivalent to an equation. Without loss of generality, and for notational conciseness, in this note we use only inequalities.
@@
Under Assumption 1, it follows from [6] that in (3) the primal is attained (i.e. the infimum is a minimum), there is no duality gap (i.e. the infimum equals the supremum) and the relaxed values are monotonically converging lower bounds on the value:
$\text{sos}(f)^{r}=\text{mom}(f)^{r}\leq\text{sos}(f)^{r+1}=\text{mom}(f)^{r+1}\leq\text{sos}(f)^{\infty}=\text{mom}(f)^{\infty}=v(f)$
where $\text{sos}(f)^{\infty}:=\lim_{r\to\infty}\text{sos}(f)^{r}$ and likewise for $\text{mom}(f)^{\infty}$.
@@ 2 Exactness cone
${\color[rgb]{0,0,0}{\mathscr{F}}({\bf g})^{r}_{d}:=\{f\in\mathbb{R}[{\bf x}]_{d}:f-v(f)\in{\mathscr{Q}}({\bf g})^{r}_{d}\}}.$
@@
Our analysis is elementary. It is inspired by the foundational work [3] which focused on the particular case of the Shor relaxation ($r=1$) with $f$ linear ($d=1$) or quadratic ($d=2$), and ${\mathscr{X}}$ a real algebraic variety defined by quadratic equations. We believe that our contribution consists of considerably simplifying and extending the geometric description of the exactness cones given in [3] to higher order relaxations of general semialgebraic sets.
@@ 3 Main result
Under Assumption 1, the exactness cone at relaxation order $r$ is given by
@@
The cone ${\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}$ is semidefinite representable since it is the projection of linear sections of the SOS cone $\Sigma[{\bf x}]$, which is itself semidefinite representable. $\Box$
@@
The exactness cone of order $r$ is semialgebraic since $f$ belongs to ${\mathscr{F}}({\bf g})^{r}_{d}$ whenever 1) $f$ belongs to ${\mathscr{Q}}({\bf g})^{r}_{d}$, a semidefinite representable hence semialgebraic cone, and 2) there exists $\hat{{\bf x}}$ in ${\mathscr{X}}$, a semialgebraic set, such that $f$ vanishes at $\hat{{\bf x}}$.
@@
It follows from the proof of Theorem 1 that it is enough to restrict the union (5) to points $\hat{{\bf x}}\in{\mathscr{X}}$ which are optimal for some objective function $f$. If $d=1$, the union can be restricted to the set of extreme points of the convex hull of ${\mathscr{X}}$, which is however not easy to describe explicitly in general. If $d\geq 2$, an anonymous reviewer noticed that every point $\hat{{\bf x}}\in{\mathscr{X}}$ is the unique global minimizer of the quadratic objective function $f({\bf x})=({\bf x}-\hat{{\bf x}})^{T}({\bf x}-\hat{{\bf x}})$, which implies that the union (5) cannot be reduced to a subset of ${\mathscr{X}}$.
@@
In our definition of the exactness cone, we require that the dual value is attained in (3). This is called SOS exactness in [1, 2]. We may relax our attainment requirements. The corresponding exactness cones would be slightly larger, at the price of more technicalities.
@@ 4 Geometry of exactness cones
Under Assumption 1, the exactness cones are monotonically embedded for increasing relaxation order:
@@
Proof: The inclusion ${\mathbb{R}}[{\bf x}]_{0}\subset{\mathscr{F}}({\bf g})^{r}_{d}$ follows from the translation invariance of the value: $v(f+a)=v(f)+a$ for all $a\in\mathbb{R}$. The inclusion ${\mathscr{F}}({\bf g})^{r}_{d}\subset{\mathscr{F}}({\bf g})^{r+1}_{d}$ follows from the inclusion relations (2). Finally, the identity $\overline{{\mathscr{F}}({\bf g})^{\infty}_{d}}=\mathbb{R}[{\bf x}]_{d}$ follows from Putinar’s Positivstellensatz – see e.g. [7, Thm. 2.15] – which states that under Assumption 1 every degree $d$ polynomial strictly positive on ${\mathscr{X}}$ belongs to ${\mathscr{Q}}({\bf g})^{r}_{d}$ for sufficiently large $r\in\mathbb{N}$, i.e. $\overline{{\mathscr{Q}}({\bf g})^{\infty}_{d}}={\mathscr{P}}({\mathscr{X}})_{d}$ and by duality $\overline{{\mathscr{R}}({\bf g})^{\infty}_{d}}={\mathscr{M}}({\mathscr{X}})_{d}$. From the definition of the exactness cone and the global minimum $\hat{v}$ on ${\mathscr{X}}$, it holds $\overline{{\mathscr{F}}({\bf g})^{\infty}_{d}}:=\{f\in\mathbb{R}[{\bf x}]_{d}:\ell_{\hat{{\bf y}}}(f)=\hat{v}=v(f),\>\hat{{\bf y}}\in{\mathscr{M}}({\mathscr{X}})_{d},\>f-\hat{v}\in{\mathscr{P}}({\mathscr{X}})_{d}\}=\mathbb{R}[{\bf x}]_{d}$. $\Box$
@@
Note that in Lemma 1 the closure is required for asymptotic exactness of all degree $d$ polynomial objective functions, i.e. $\overline{{\mathscr{F}}({\bf g})^{\infty}_{d}}=\mathbb{R}[{\bf x}]_{d}$. A classical example is $n=m=1$, $g_{1}(x)=-x^{2}$ for which ${\mathscr{X}}=\{0\}$. Whereras $f(x)=\pm x\notin{\mathscr{Q}}({\bf g})^{r}_{1}$ for finite $r\in\mathbb{N}$, it holds that $f(x)+\varepsilon=\frac{\varepsilon}{2}+\frac{\varepsilon}{2}(1\pm\frac{x}{\varepsilon})^{2}-\frac{x^{2}}{2\varepsilon}\in{\mathscr{Q}}({\bf g})^{{\color[rgb]{0,0,0}2}}_{1}$ for every $\varepsilon>0$, see [1, Ex. 1.3.4] or [8, §2.5.2]. For this example ${\mathscr{F}}({\bf g})^{r}_{1}=\mathbb{R}[{\bf x}]_{0}$ for all finite $r$, and $\overline{{\mathscr{F}}({\bf g})^{\infty}_{1}}=\mathbb{R}[{\bf x}]_{1}$.
The vector space of polynomials $\mathbb{R}[{\bf x}]_{d}$ can be identified with ($\mathbb{R}^{n+d\choose n}$. Using the linear functional $\ell_{{\bf y}}(f)$, it is in duality with the space of vectors ${\bf y}\in\mathbb{R}^{n+d\choose n}$. Given $\hat{{\bf x}}\in{\mathscr{X}}$, let ${\bf y}_{\hat{{\bf x}}}{\>\in\>\mathbb{R}^{n+d\choose n}}$ be the Dirac vector at $\hat{{\bf x}}$, i.e. such that $\ell_{{\bf y}_{\hat{{\bf x}}}}(f)=f(\hat{{\bf x}})$ for every $f\in\mathbb{R}[{\bf x}]_{d}$. Given a convex set ${\mathscr{Y}}{\>\subset\>}\mathbb{R}^{N}$, let ${\mathscr{N}}_{{\mathscr{Y}}}(\hat{{\bf y}}):=\{f\in\mathbb{R}[{\bf x}]_{d}:\ell_{{\bf y}-\hat{{\bf y}}}(f)\leq 0,\>\forall\>{\bf y}\in{\mathscr{Y}}\}$ denote the normal cone to ${\mathscr{Y}}$ at a point $\hat{{\bf y}}\in{\mathscr{Y}}$. With these notations, we have the following geometric counterpart to Theorem 1.
Theorem 2.
Up to the sign and closure, the spectrahedral cones of Theorem 1 are normal cones to the moment relaxation:
${\mathscr{N}}_{{\mathscr{R}}({\bf g})^{r}_{d}}({\bf y}_{\hat{{\bf x}}})={\color[rgb]{0,0,0}-\overline{{\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}}}$
@@
Proof: From the definition of the normal cone: $-{\mathscr{N}}_{{\mathscr{R}}({\bf g})^{r}_{d}}({\bf y}_{\hat{{\bf x}}})=\{f\in\mathbb{R}[{\bf x}]_{d}:\ell_{{\bf y}}(f)\geq\ell_{{\bf y}_{\hat{{\bf x}}}}(f)=f(\hat{{\bf x}})\>\>\forall{\bf y}\in{\mathscr{R}}({\bf g})^{r}_{d}\}=\{f\in\mathbb{R}[{\bf x}]_{d}:\ell_{{\bf y}}(f-f(\hat{{\bf x}}))\geq 0\>\>\forall{\bf y}\in{\mathscr{R}}({\bf g})^{r}_{d}\}$ which is the same as $\{f\in\mathbb{R}[{\bf x}]_{d}:f-f(\hat{{\bf x}})\in{\color[rgb]{0,0,0}\overline{{\mathscr{Q}}({\bf g})^{r}_{d}}}\}={\color[rgb]{0,0,0}\overline{{\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}}}$ since $\overline{{\mathscr{Q}}({\bf g})^{r}_{d}}$ is the dual cone to ${\mathscr{R}}({\bf g})^{r}_{d}$. $\Box$
Theorem 2 is a generalization of [3, Prop. 4.7] which considers only the case $d\leq 2$ and the Shor relaxation (i.e. $r=1$) for real varieties ${\mathscr{X}}$ generated by quadratic polynomials.
If ${\mathscr{X}}$ has an interior point, then the truncated quadratic module ${\mathscr{Q}}({\bf g})^{r}_{d}$ is closed, and the dual is attained in (3) (i.e. the supremum is a maximum). If ${\mathscr{X}}$ does not have an interior point, e.g. if it is a low-dimensional algebraic variety, then additional algebraic or geometric conditions are required for the dual to be attained, cf. [2, 1, 8]. This is why the closure is required in the statement of Theorem 2.
It is of interest to know whether we have exactness at relaxation order $r$ for all objective functions. A sufficient condition is that ${\mathscr{R}}({\bf g})^{r}_{d}={\mathscr{M}}({\mathscr{X}})_{d}$, i.e. all vectors of the moment relaxation are convex combinations of Dirac vectors of ${\mathscr{X}}$. If $d=1$ the condition becomes ${\mathscr{R}}({\bf g})^{r}_{1}=\mathbb{R}_{+}\times\mathrm{conv}\>{\mathscr{X}}$, the Cartesian product of the positive real line with the convex hull of ${\mathscr{X}}$.